# The Rate-Distortion-Perception-Classification Tradeoff: Joint Source Coding and Modulation via Inverse-Domain GANs

### v1

@@ II-B Tradeoff Analyses
The study of tradeoffs in lossy compression can be traced back to rate-distortion theory [1], which characterizes the rate-distortion function
$R(D)=\begin{array}[t]{cl}\underset{p_{\widehat{\bm{X}}|\bm{X}}}{\min}&I(\widehat{\bm{X}},\bm{X})\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\widehat{\bm{X}},\bm{X})\big{]}\leq D\,,\end{array}$ (10)
@@
where $I(\widehat{\bm{X}},\bm{X})$ is the mutual information between $\bm{X}$ and its reconstruction $\widehat{\bm{X}}$. The $R(D)$ function has a closed-form expression under some simple source distributions and distortion metrics. Recent work has gone beyond using reconstruction metrics, e.g., the mean squared error (MSE), to assess image quality, considering also perception and semantic metrics.
@@
The PD tradeoff. For example, [3] studied the perception-distortion (PD) tradeoff by replacing the objective in (10) with a divergence metric $d(p_{\bm{X}},p_{\widehat{\bm{X}}})$ [cf. (7)]. Assuming that the input signal follows a Rademacher distribution, they proved the existence of a tradeoff between the best achievable divergence and the allowable distortion $D$.
@@
The RDP tradeoff. Building on [3], [4] studied the rate-distortion-perception (RDP) tradeoff. The problem they analyzed was a variation of (7), without the last constraint (on classification error) and with $I(\widehat{\bm{X}},\bm{X})$ in the objective, instead of the rate. Assuming a Bernoulli input, they showed that in lossy image compression, the higher the perception quality of the output images, the lower the achievable rate. Although insightful, the analysis in [4] is not applicable to our scenario, as it considers only scalar signals, thus ignoring the possibility of compressing them, and also skips the quantization step. The work in [25] further improved on [4] and showed that, for a fixed bit rate, imposing a perfect perception constraint doubles the lowest achievable MSE. It further proposed a training framework to achieve the lowest MSE distortion under a perfect perception constraint at a given bit rate.
@@
The CDP tradeoff. The work in [5] analyzed instead the classification-distortion-perception (CDP) tradeoff, i.e., a modification of problem (7) in which $\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})$ is minimized under the first two constraints (and the rate is ignored). Assuming an input signal that is drawn from a Gaussian mixture model with two classes, they showed the existence of a tradeoff. Our setup is more general, as we do not require the input to be Gaussian nor to be drawn from just two classes.
@@
Our approach. In all the above work, the signals are assumed scalar, which is not suitable to study compression in terms of dimensionality reduction. By contrast, in (7), we consider vector signals and minimize the channel rate subject to constraints in distortion, perception, and classification error. Furthermore, we show the existence of a strict tradeoff, rather than just a simple tradeoff (as in [3, 4, 5]) between rate and all the constraints of (7).
@@ III The RDPC Tradeoff
Recall our multiclass signal model in (1) and the channel model in (2). Recall also that we assume a Gaussian channel $p_{\widehat{\bm{Y}}|\bm{Y}}(\widehat{\bm{y}}|\bm{y})=\mathcal{N}\big{(}\bm{y}\,,\bm{\Sigma}\big{)}$, where $\bm{\Sigma}_{ii}>0$ is the noise power of channel $i$, and $\bm{\Sigma}_{ij}=0$ for $i\neq j$.
We assume a deterministic classifier $c_{0}\,\mathrel{\mathop{\ordinarycolon}}\,\mathbb{R}^{n}\to\{1,\ldots,L\}$ which, for $l=1,\ldots,L$, decides $c_{0}(\widehat{\bm{X}})=l$ whenever $\widehat{\bm{X}}$ belongs to a fixed region $\mathcal{R}_{l}\subset\mathbb{R}^{n}$. The probability of classification error is then
$\epsilon_{c_{0}}(\bm{X},\,\widehat{\bm{X}})$ $=\mathbb{P}\big{(}\text{class}(\bm{X})\neq c_{0}\big{(}\widehat{\bm{X}}\big{)}\big{)}$
@@
$=\sum_{i<j}p_{j}\cdot\int_{\mathcal{R}_{i}}\operatorname{d\!}\,p_{\widehat{\bm{X}}|H_{j}}\,,$ (11)
where $p_{j}\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}(H_{l})$ is the probability of $\bm{X}$ being drawn from class $j$. Our main result is as follows.
@@
Theorem 1 is generic and applies to any distortion metric $\Delta$, perception metric $d$, and classifier $c_{0}$. The main assumption is that the perception metric $d(\cdot,\cdot)$ is convex in the second argument, which holds for a variety of divergences, e.g., $f$-divergence (including total variation, Kullback-Leibler, and Hellinger distance) and Rényi divergence [26, 27]. The same assumption was used in [3, 4, 5]. The theorem says that if we optimize the channel for the smallest possible rate, the encoding-decoding system cannot achieve arbitrarily small distortion, semantic error, and classification error. These metrics are in conflict and we need to strike a tradeoff between them. This behavior will be observed in practice when we design algorithms to (approximately) solve the RDPC problem in (7). Note that while prior work [3, 4, 5] shows the existence of a tradeoff by proving that a certain function is convex in each argument, we establish a strict tradeoff by proving that $R(D,P,C)$ is strictly convex in each argument.
IV Algorithms for Joint Source Coding and Modulation
As solving the RDPC problem in (7) in full generality (i.e., non-parametrically) is difficult, we propose two algorithms that approximately solve that problem. Our main algorithm is described in Section IV-A and is inspired on inverse-domain GAN (ID-GAN) [7]. As will be shown in Section V, it can handle extreme compression ratios while preserving distortion, perception, and classification metrics. Then, in Section IV-B, we propose an algorithm that tries to directly solve (7) under a set of simplifying assumptions. This will enable to draw some insights on the performance of the former algorithm.
IV-A Inverse-domain GAN compression
As formulated in (7), our goal is to design an encoder-decoder pair and channel noise levels that preserve image reconstruction quality, perception quality, and classification accuracy. Neural networks, and GANs [6] in particular, are a powerful tool to achieve this. In our channel diagram (2), we will thus model the encoder $p_{\bm{Y}|\bm{X}}$ with a neural network $e(\cdot\,;\bm{\theta}_{e})\,\mathrel{\mathop{\ordinarycolon}}\,\mathbb{R}^{n}\to\mathbb{R}^{m}$ parameterized by $\bm{\theta}_{e}$, and the decoder $p_{\widehat{\bm{X}}|\widehat{\bm{Y}}}$ as neural network $d(\cdot\,;\bm{\theta}_{d})\,\mathrel{\mathop{\ordinarycolon}}\,\mathbb{R}^{m}\to\mathbb{R}^{n}$ parameterized by $\bm{\theta}_{d}$. These networks will be trained as in ID-GAN [7] which, however, was proposed for a task different from JSCM. Specifically, given an (adversarially-trained) image generator, the goal in [7] was to train an encoder to obtain a semantically-meaningful latent code for image editing. We adopt this process of training the generator first, and then the encoder.
[…]
Figure 1: Proposed ID-GAN framework for solving the RDPC problem in JSCM. The generator/decoder is first trained adversarially with critic 1 in the first step. The decoder is then fixed and coupled with an encoder, which is in turn trained with critic 2 in order to preserve both reconstruction quality and classification accuracy. Critics 1 and 2 have the same architecture.
Proposed scheme. Fig. 1 shows our framework for JSCM. As in ID-GAN [7], we first train an image generator/decoder $d(\cdot\,;\,\bm{\theta}_{d})$ (Fig. 1, top) adversarially against discriminator $f_{1}$, which learns to distinguish a real signal from a randomly generated one, $d(\bm{Z}\,;\,\bm{\theta}_{d})$, where $\bm{Z}\sim\mathcal{N}(\bm{0_{m}},\,\bm{I_{m}})$ is a vector of i.i.d. standard Gaussians. This is the conventional GAN setup [6, 28]. As the discriminator is a particular case of a classifier, outputting just a binary signal, it is also known as a critic. Once the decoder is trained, we fix it and train the encoder $e(\cdot\,;\,\bm{\theta}_{e})$ together with its own critic $f_{2}$, which again learns to distinguish real signals from randomly generated from ones (Fig. 1, bottom). This will enable us to compute semantically-meaningful latent codes to be transmitted along a communication channel. In other words, in (2), $p_{\bm{Y}|\bm{X}}$ is implemented by $e(\cdot\,;\,\bm{\theta}_{e})$, and $p_{\widehat{\bm{Y}}|\widehat{\bm{X}}}$ is implemented by $d(\cdot\,;\,\bm{\theta}_{d})$. To apply the scheme in a JSCM scenario, we normalize the power in the last layer of encoder and evaluate the system for different transmission rates. Next, we explain the training processes of the decoder and encoder in more detail.
Training the decoder. To train the decoder $d(\cdot\,;\,\bm{\theta}_{d})$ and critic $f_{1}$ as in Fig. 1 (top), we leverage the Wasserstein GAN (WGAN) [28] framework. This consists of finding the parameters $\bm{\theta}_{d}$ of the decoder that minimize the Wasserstein-1 (or earth-mover) distance $W_{1}(p_{r},p_{\bm{\theta}_{d}})$ between the distribution $p_{r}$ of real data and the distribution $p_{\bm{\theta}_{d}}$ of data generated by $d(\bm{Z}\,;\,\bm{\theta}_{d})$, with $\bm{Z}\sim\mathcal{N}(\bm{0},\,\bm{I})$:
where the supremum is over the functions $f_{1}\,\mathrel{\mathop{\ordinarycolon}}\,\mathbb{R}^{n}\to\mathbb{R}$ that are $1$-Lipschitz continuous. Notice that (12) is given in Kantarovich-Rubinstein dual form. [^4] The definition of the Wasserstein-$p$ distance between two probability measures $p_{\bm{X}},p_{\bm{Y}}$ in $\mathbb{R}^{n}$ is $W_{p}(p_{\bm{X}},p_{\bm{Y}})=\big{(}\underset{\gamma\in\Pi(p_{\bm{X}},p_{\bm{Y}})}{\inf}\mathbb{E}_{(\bm{X},\bm{Y})\sim\gamma}\big{[}\|\bm{X}-\bm{Y}\|_{2}^{p}\big{]}\big{)}^{\frac{1}{p}}$, where $\Pi\big{(}p_{\bm{X}},p_{\bm{Y}}\big{)}$ is the set of all joint distributions with marginals $p_{\bm{X}}$ and $p_{\bm{Y}}$, and $1\leq p\leq+\infty$. While the critic $f_{1}$ is found by maximizing the argument in (12), the parameters of the decoder are found by minimizing the full Wasserstein distance $W_{1}(p_{r},p_{\bm{\theta}_{d}})$. This distance enables overcoming the mode collapse observed in the original GAN framework [6], which used instead the Jensen-Shannon divergence. To enforce Lipschitz-continuity of the critic $f_{1}$, [28] proposed to limit its parameters to a small box around the origin. The work in [29], however, found that this technique leads to instabilities in training (exploding/vanishing gradients) and showed that a gradient penalty solves these problems. We thus adopt the loss suggested in [29] for finding critic $f_{1}$:
where $\lambda_{g}\geq 0$, and $\widetilde{\bm{X}}$ is a point sampled uniformly in the line joining a real data point $\bm{X}\sim p_{r}$ and a point $\bm{Y}\sim p_{\bm{\theta}_{d}}$ generated by the decoder. Notice that the third term in (13) eliminates the need to constrain the critic to be Lipschitz-continuous [constraint in (12)]; see [29] for more details. In turn, the parameters $\bm{\theta}_{d}$ of the decoder are found by minimizing (12) which, when $f_{1}$ is fixed, is equivalent to minimizing
$L_{\bm{\theta}_{d}}=-\mathbb{E}_{\bm{X}\sim p_{\bm{\theta_{d}}}}[f_{1}(\bm{X})]=-\mathbb{E}_{\bm{Z}\sim\mathcal{N}(\bm{0},\,\bm{I})}\big{[}f_{1}\big{(}d(\bm{Z}\,;\,\bm{\theta}_{d})\big{)}\big{]}\,.$ (14)
During training, we update the parameters of the critic $n_{\text{critic}}$ times while keeping the parameters of the decoder fixed. This guarantees that the supremum in (12) is reasonably well computed. See [28, 29] for details.
Algorithm 1 ID-GAN compression: training of the encoder
1:Training images/labels $\{(\bm{x}^{(t)},\bm{\ell}^{(t)})\}_{t=1}^{T}$, pretrained decoder $d(\cdot)$, pretrained classifier $c(\cdot)$, learning rate $\alpha$, momentum parameters $\beta_{1}$, $\beta_{2}$, batch size $S$, number of iterations of critic $n_{\text{critic}}$, loss hyperparameters $\lambda_{g}$, $\lambda_{d}$, $\lambda_{p}$, $\lambda_{c}$
2:Set encoder $\bm{\theta}_{e}^{(0)}$ and critic $\bm{\theta}_{f}^{(0)}$ parameters randomly
3:
4:
5:$\mathcal{S}=\text{randperm}(\{1,2,\ldots,T\})$
6:for $j=1,\ldots,\lceil T/S\rceil$ do
7:    Select next $S$ batch indices $\mathcal{S}_{j}$ from $\mathcal{S}$
8:    for $k=1$ to $n_{\text{critic}}$ do
9:         $L_{f_{2}}^{(1)}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}f_{2}\Big{(}d\big{(}e(\bm{x}^{(s)}\,;\,\bm{\theta}_{e}^{(j)})\big{)}\,;\,\bm{\theta}_{f}^{(k)}\Big{)}$
10:         $L_{f_{2}}^{(2)}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}f_{2}\Big{(}\bm{x}^{(s)}\,;\,\bm{\theta}_{f}^{(k)}\Big{)}$
11:         for $s\in\mathcal{S}_{j}$ do
12:             Draw $\epsilon\sim\mathcal{U}(0,1)$ randomly
13:             $\widetilde{\bm{x}}^{(s)}=(1-\epsilon)\cdot\bm{x}^{(s)}+\epsilon\cdot d\big{(}e\big{(}\bm{x}^{(s)}\,;\,\bm{\theta}_{e}^{(j)}\big{)}\big{)}$
14:         end for
15:         $L_{f_{2}}^{(3)}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}\Big{(}\big{\|}\nabla_{\widetilde{\bm{x}}}f_{2}(\widetilde{\bm{x}}^{(s)})\big{\|}_{2}-1\Big{)}^{2}$
16:         $L_{f_{2}}=L_{f_{2}}^{(1)}-L_{f_{2}}^{(2)}+\lambda_{g}L_{f_{2}}^{(3)}$
17:         $\bm{\theta}_{f}^{(k+1)}=\text{Adam}\Big{(}\bm{\theta}_{f}^{(k)},\,L_{f_{2}},\,\alpha,\,\beta_{1},\,\beta_{2}\Big{)}$
18:    end for
19:    $\bm{\theta}_{f}^{(j)}=\bm{\theta}_{f}^{(n_{\text{critic}})}$
20:    for $s\in\mathcal{S}_{j}$ do
21:         $\widehat{\bm{x}}^{(s)}=d\Big{(}e\big{(}\bm{x}^{(s)}\,;\,\bm{\theta}_{e}^{(j)}\big{)}\Big{)}$
22:    end for
23:
$L_{e}=\frac{1}{S}\sum_{s\in\mathcal{S}_{j}}-f_{2}\Big{(}\widehat{\bm{x}}^{(s)}\,;\,\bm{\theta}_{f}^{(j)}\Big{)}+\lambda_{d}\Big{\|}\bm{x}^{(s)}-\widehat{\bm{x}}^{(s)}\Big{\|}_{2}^{2}\\[-5.69046pt] +\lambda_{c}\text{CE}\big{(}c(\widehat{\bm{x}}^{(s)}),\bm{\ell}^{(s)}\big{)}$
24:    $\bm{\theta}_{e}^{(j+1)}=\text{Adam}\Big{(}\bm{\theta}_{e}^{(j)},\,L_{e},\,\alpha,\,\beta_{1},\,\beta_{2}\Big{)}$
25:end for
Training the encoder. After training the decoder $d(\cdot\,;\,\bm{\theta}_{d})$, we fix it and consider the scheme in Fig. 1 (bottom) to train the encoder $e(\cdot\,;\,\bm{\theta}_{e})$. As the decoder, the encoder is also trained adversarially against a critic $f_{2}$, but now we take into account the quality of reconstruction and its semantic meaning. The former is captured by an MSE loss between the original and reconstructed images, and the latter by a cross-entropy loss between the image label and the output of a pre-trained classifier $c(\cdot)$ applied to the reconstructed image. The full loss is associated to the encoder is
$L_{e}=-\mathbb{E}\big{[}f_{2}(d(e(\bm{X})))\big{]}+\mathbb{E}\Big{[}\lambda_{d}\big{\|}\bm{X}-\widehat{\bm{X}}\big{\|}_{2}^{2}\\ +\lambda_{c}\text{CE}\Big{(}c\big{(}\widehat{\bm{X}}\big{)},\text{class}\big{(}\bm{X}\big{)}\Big{)}\Big{]}\,,$ (15)
[…]
$L_{f_{2}}=\mathbb{E}\Big{[}f_{2}\big{(}d(e(\bm{X}))\big{)}\Big{]}-\mathbb{E}\big{[}f_{2}(\bm{X})\big{]}\\ +\lambda_{g}\,\mathbb{E}_{\bm{X},\epsilon}\Big{[}\big{(}\big{\|}\nabla_{\widetilde{\bm{x}}}f_{2}\big{(}\widetilde{\bm{X}}\big{)}\big{\|}_{2}-1\big{)}^{2}\Big{]}\,,$ (16)
where $\widetilde{\bm{X}}=(1-\epsilon)\bm{X}+\epsilon\,d(e(\bm{X}))$ and $\epsilon\sim\mathcal{U}(0,1)$ is uniformly distributed in $[0,1]$. Note that we omitted the parameters of the encoder and decoder for simplicity.
The complete training procedure of the encoder is shown in Algorithm 1. Its inputs include training images $\bm{x}^{(t)}$ and corresponding labels $\bm{\ell}^{(t)}$, and a pretrained decoder $d(\cdot)$ and classifier $c(\cdot)$. After initializing the parameters of the encoder and associated critic, in each epoch we randomly permute the indices of the training data (step 5) and visit all the training data in batches of size $S$. This takes $\lceil T/S\rceil$ iterations, where $T$ is the number of data points. The loop in steps 8-18 performs $n_{\text{critic}}$ iterations of Adam to minimize the critic loss in (16) and thus to update the critic $f_{2}$ parameters $\bm{\theta}_{f}$ (where we omit the index $2$ for simplicity). This corresponds to computing the supremum in the Wasserstein distance (12) between the real data $\bm{X}$ and the reconstructed one $\widehat{\bm{X}}=d\big{(}e\big{(}\bm{X}\big{)}\big{)}$. The terms in (16) are computed separately, with the last term requiring the creation of the intermediate variables $\widetilde{\bm{x}}^{(s)}$ in steps 11-14. As usual, expected values were replaced by sample averages over the batch. After having updated the parameters of critic $f_{2}$, we perform one iteration of Adam to minimize the encoder loss in (15) and thus to update the encoder parameters $\bm{\theta}_{e}$. This requires passing each image in the batch through the encoder and decoder to create $\widehat{\bm{x}}^{(t)}$, as in steps 20-22.
Connection with (7). Although minimizing (15) constitutes a problem different from (7), the three terms in (15) capture the quantities constrained in (7). That is, $\mathbb{E}\big{[}\|\bm{X}-\widehat{\bm{X}}\|_{2}^{2}\big{]}$ in (15) is a particular instance of $\mathbb{E}[\Delta(\bm{X},\widehat{\bm{X}})]$ in the distortion constraint of (7); the term $L_{f_{2}}$ in (15) approximates the Wasserstein distance between the distributions of $\bm{X}$ and $\widehat{\bm{X}}$, a particular instance of $d(p_{\bm{X}},p_{\widehat{\bm{X}}})$ in the perception constraint of (7); and the cross-entropy term in (15) is a proxy of the classification error term in the last constraint of (7). However, rather than minimizing the rate as in (7), here we fix the dimension $m$ of the latent variable and evaluate the performance of the resulting scheme for different values of $m$. Next, we attempt to solve (7) directly under some simplifying assumptions.
IV-B Solving RDPC: Problem Formulation
To develop insight on how the proposed framework strikes a tradeoff between achievable rate, distortion, perception, and classification error, we now analyze a simplified version of problem (7). We make the following assumptions.

### v2

@@ II-C Tradeoff analyses
The study of tradeoffs in lossy compression can be traced back to rate-distortion theory [cover1999elements], which characterizes the rate-distortion function
$R(D)=\begin{array}[t]{cl}\underset{p_{\widehat{\bm{X}}|\bm{X}}}{\min}&I(\bm{X},\widehat{\bm{X}})\\ \text{s.t.}&\mathbb{E}\big{[}\Delta(\bm{X},\widehat{\bm{X}})\big{]}\leq D\,,\end{array}$ (10)
@@
where $I(\bm{X},\widehat{\bm{X}})$ is the mutual information between $\bm{X}$ and its reconstruction $\widehat{\bm{X}}$. The $R(D)$ function has a closed-form expression under some simple source distributions and distortion metrics. Recent work has gone beyond using reconstruction metrics, e.g., the mean squared error (MSE), to assess image quality, considering also perception and semantic metrics.
@@
The PD tradeoff. For example, [blau2018perception] studied the perception-distortion (PD) tradeoff by replacing the objective in (10) with a divergence metric $d(p_{\bm{X}},p_{\widehat{\bm{X}}})$ [cf. (7)]. Assuming that the input signal follows a Rademacher distribution, they proved the existence of a tradeoff between the best achievable divergence and the allowable distortion $D$.
@@
The RDP tradeoff. Building on [blau2018perception], [RDPtradeoff] studied the rate-distortion-perception (RDP) tradeoff. The problem they analyzed was a variation of (7), without the last constraint (on classification error) and with $I(\bm{X},\widehat{\bm{X}})$ in the objective, instead of the rate. Assuming a Bernoulli input, they showed that in lossy image compression, the higher the perception quality of the output images, the lower the achievable rate. Although insightful, the analysis in [RDPtradeoff] is not applicable to our scenario, as it considers only scalar signals, thus ignoring the possibility of compressing them, and also skips the quantization step. The work in [yan2021perceptual] further improved on [RDPtradeoff] and showed that, for a fixed bit rate, imposing a perfect perception constraint doubles the lowest achievable MSE. It further proposed a training framework to achieve the lowest MSE distortion under a perfect perception constraint at a given bit rate.
@@
The CDP tradeoff. The work in [CDP] analyzed instead the classification-distortion-perception (CDP) tradeoff, i.e., a modification of problem (7) in which $\mathbb{E}[\epsilon_{c_{0}}(\bm{X},\widehat{\bm{X}})]$ is minimized subject to the first two constraints (the rate is ignored). Assuming an input signal that is drawn from a Gaussian mixture model with two classes, they showed the existence of a tradeoff. Our setup is more general, as we do not require the input to be Gaussian nor to be drawn from just two classes.
@@
Our approach. In all the above work, the signals are assumed scalar, which is not suitable to study compression in terms of dimensionality reduction. By contrast, in (7), we consider vector signals and minimize the channel rate subject to constraints on distortion, perception, and classification error. Furthermore, we show the existence of a strict tradeoff, rather than just a simple tradeoff (as in [blau2018perception, RDPtradeoff, CDP]) between rate and all the constraints of (7).
@@ III The RDPC Tradeoff
We now establish the existence of an inherent tradeoff in solving problem (7). Recall our multiclass signal model in (1) and the channel model in (2). Recall also that we assume a Gaussian channel $p_{\widehat{\bm{Y}}|\bm{Y}}(\widehat{\bm{y}}|\bm{y})=\mathcal{N}\big{(}\bm{y}\,,\bm{\Sigma}\big{)}$, where $\bm{\Sigma}_{ii}>0$ is the noise power of channel $i$, and $\bm{\Sigma}_{ij}=0$ for $i\neq j$.
We assume a deterministic classifier $c_{0}\,:\,\mathbb{R}^{n}\to\{1,\ldots,L\}$ which, for $l=1,\ldots,L$, decides $c_{0}(\widehat{\bm{X}})=l$ whenever $\widehat{\bm{X}}$ belongs to a fixed region $\mathcal{R}_{l}\subset\mathbb{R}^{n}$. Assuming $\epsilon_{c_{0}}$ is the 00-$1$ loss, the expected classification error is then
$\mathbb{E}\Big{[}\epsilon_{c_{0}}(\bm{X},\,\widehat{\bm{X}})\Big{]}$ $=\mathbb{P}\big{(}\text{class}(\bm{X})\neq c_{0}\big{(}\widehat{\bm{X}}\big{)}\big{)}$
@@
[…]
where $p_{j}:=\mathbb{P}(H_{l})$ is the probability of $\bm{X}$ being drawn from class $j$. Our main result is as follows.
@@
Theorem 1 is generic and applies to any distortion metric $\Delta$, perception metric $d$, and classifier $c_{0}$. The main assumption is that the perception metric $d(\cdot,\cdot)$ is convex in the second argument, which holds for a variety of divergences, e.g., $f$-divergence (including total variation, Kullback-Leibler, and Hellinger distance) and Rényi divergence [Csiszar04-InformationTheoryAndStatistics, Erven14-RenyiKL]. The same assumption was used in [blau2018perception, RDPtradeoff, CDP]. The theorem says that if we optimize the channel for the smallest possible rate, the encoding-decoding system cannot achieve arbitrarily small distortion, perception error, and classification error. These metrics are in conflict and we need to strike a tradeoff between them. This behavior will be observed in practice when we design algorithms to (approximately) solve the RDPC problem. Note that while prior work [blau2018perception, RDPtradeoff, CDP] shows the existence of a tradeoff by proving that a certain function is convex in each argument, we establish a strict tradeoff by proving that $R(D,P,C)$ is strictly convex in each argument.
As solving the RDPC problem in (7) in full generality i.e., non-parametrically, is difficult, in the next two sections we propose two algorithms that approximately solve that problem under different assumptions. As we will see in the experiments in Section VI, both algorithms exhibit the tradeoff behavior stipulated by Theorem 1.
IV RDPC Tradeoff under GMM signals and linear encoder and decoder
To make problem (7) more tractable, in this section we assume that the source signal $\bm{X}$ in (1) is a Gaussian mixture model (GMM) drawn from two classes and that the encoder and decoder are linear. This will enable us to approximate (7) with a problem whose optimal cost function upper bounds the optimal cost of (7) (Section IV-A). We then develop an algorithm, RDPCO, to solve the resulting problem (Section IV-B). More formally, we make the following assumptions.